# High-cadence observations of galactic nuclei by the future two-band UV-photometry mission QUVIK

### v1

@@ 3.3 Repeating nuclear transients
Repeating nuclear transients have been associated with stars or compact remnants orbiting SMBHs, causing recurrent flares on the timescales of several months to hours (see Suková et al.[124] for a review). Repetitive outbursts on longer timescales are likely caused by stripping the outer layers of a star while the inner core remains intact (partial TDEs). This is the case when the relative impact parameter $\beta=r_{\rm t}/r_{\rm p}<\beta_{\rm crit}$, where $r_{\rm t}$ is the tidal radius and $r_{\rm p}$ is the pericenter distance of the star. For main-sequence stars, $\beta_{\rm crit}$ was constrained to be $\sim 2-3$[125] and for these values, the star likely gets fully disrupted. Partial TDEs exhibit variability in both UV/optical and X-ray wavebands, often with a time delay due to different processes (e.g. AT2018fyk[126] or ASASSN0-14ko[127]). The shortest repeating nuclear transients are Quasiperiodic Erupters (QPEs), which exhibit rapid large-amplitude soft X-ray flares on the timescale of hours to days; however, they do not exhibit significant variability in the optical/UV domains. Recently, a class of repeating soft X-ray nuclear transients on the intermediate timescales of tens of days has been revealed (represented by Swift J0230-28[18, 19]) and they likely correspond to partial TDEs of non-main-sequence stars, such as red giants or even gas-giant planets[20]. In the case of Swift J0230-28, no significant variability in the optical/UV bands was found, though a radio flare was associated with the start of one of the X-ray flares[19].
@@
QPEs[128, 129] are a class of fast, quasiperiodic large-amplitude outbursts observed in the nuclei of a few galaxies. These outbursts exhibit recurrence times ranging from hours to days and have been detected primarily in the X-ray band, with peak luminosities of $\sim 10^{42}$ erg s-1. The exact mechanism for their production is uncertain – the proposed models range from accretion-disk instabilities[130, 131] to bodies periodically crossing an accretion disk[132, 124]. A favored class of models involve extreme mass-ratio inspirals (EMRIs)[133], i.e. stars or compact remnants which collide with the accretion disks of supermassive black holes (SMBHs), producing flares through disk–star interactions involving shocks within the ejected material [134]. This class of models can also address the variability in the flare recurrence timescale, in particular the long-short recurrence timescale pattern, which can be attributed to the EMRI orbital precession in combination with the accretion-disk precession [134, 135].
@@
QPEs have been detected in the soft X-ray domain so far, with no significant variability in the optical and UV wavebands. However, since the EMRI distance can be expressed in gravitational radii as $r_{\rm EMRI}/r_{\rm g}\propto M_{\bullet}^{-2/3}P_{\rm QPE}^{2/3}$, longer-period QPEs are expected to result in softer flares peaking in the UV domain for the fixed SMBH mass[136]. Similarly, for the QPE periodicity comparable to the X-ray QPEs, the flares are expected to be softer and potentially detectable in the UV domain in lower-mass SMBH systems. Also, since some X-ray QPEs seem to be associated with post-TDE nuclei with a decreasing relative accretion rate, the flare spectra are expected to get softer with the drop in the accretion rate, potentially ceasing to be detected in the X-ray domain and appearing in the UV bands [136] when the relative accretion rate reaches $\dot{m}\sim\mathcal{O}(10^{-2})$. The detection of such events in the UV domain is feasible with future missions such as ULTRASAT and QUVIK, which have sensitivities down to $\sim 22$ AB magnitude. The QUVIK mission has the projected sensitivity limit of $\sim 22$ AB mag in the NUV band (325 nm), which implies the monochromatic flux density limit of $F_{\nu}\sim 5.75\times 10^{-29}\,{\rm erg\,s^{-1}\,cm^{-2}\,Hz^{-1}}$ or $\nu F_{\nu}\sim 5.31\times 10^{-14}\,{\rm erg\,s^{-1}\,cm^{-2}}$ in the NUV band. UV QPEs in galaxies could thus be detectable within a distance of up to $d_{\rm L}=235(L_{\rm QPE}/3.5\times 10^{41}\,{\rm erg\,s^{-1}})^{1/2}(\nu F_{\nu}/5.31\times 10^{-14}\,{\rm erg\,s^{-1}\,cm^{-2}})^{-1/2}\,{\rm Mpc}$ ($z=0.053$), assuming the typical UV QPE luminosity of $3.5\times 10^{41}\,{\rm erg\,s^{-1}}$[136]. The detection and monitoring of UV QPEs would significantly advance our understanding of repeating nuclear transients, their potential association with EMRIs, and the evolution of SMBH accretion disks plausibly following a TDE[137].
4 Synergy with other missions
@@
    the telescope will perform dedicated monitoring of the variability of selected bright AGN ($m_{\rm AB}\lesssim 18$ mag) located at low to intermediate redshifts ($z\lesssim 1$). These AGN programs and campaigns will be selected based on a competitive peer-review process,
@@ 5 Summary
QUVIK is an approved versatile two-band UV photometry space mission currently under development. It is designed to perform observations of the early UV emission of kilonovae, which is its primary scientific objective. In addition, it will conduct monitoring and follow-up observations of various sources, including stellar transients, hot stars, binary stars, stellar associations and clusters. Building on the legacy of previous successful UV space telescopes, QUVIK will also bring valuable results concerning active galactic nuclei (AGN) and nuclear transients. The main areas of the AGN research include:
  • reverberation mapping of accretion disks. QUVIK can conduct high-cadence observations of AGN variability simultaneously in both FUV and NUV bands. For the sources with the SMBH mass of $10^{9}\,M_{\odot}$, the required cadence is at least $\sim 1$ day, while for $10^{8}\,M_{\odot}$, the cadence needs to increase to $\sim 0.2$ days,
@@
QUVIK will operate in synergy with wide field-of-view observing facilities, including ULTRASAT, Vera C. Rubin Legacy Survey of Space and Time as well as with gravitational-wave observatories (LIGO-Virgo-KAGRA). In addition, the legacy of QUVIK will prepare the path for future UV missions, such as UVEX. Its potential detections of repeating nuclear transients, some of which may be associated with extreme- and intermediate-mass ratio inspirals, will be beneficial for the upcoming space-borne low-frequency gravitational-wave observatories (LISA).
@@ Acknowledgments
MZ, HB, MP, VK, MM, and PS acknowledge the support from the Czech Science Foundation (GACR) Junior Star grant no. GM24-10599M. NW and ILG were supported by the Czech Science Foundation (GACR) grant 21-13491X. PK and ML acknowledge the continuing support of the OPUS-LAP/GAČR-LA bilateral research grant (2021/43/I/ST9/01352/OPUS22 and GF23-04053L). JL acknowledges the support from the J.W. Fulbright Commission in the Czech Republic.

### v2

@@ 4.3 Repeating nuclear transients
Repeating nuclear transients have been associated with stars or compact remnants orbiting SMBHs, causing recurrent flares on the timescales of several months to hours (see Suková et al.[139] for a review). Repetitive outbursts on longer timescales are likely caused by stripping the outer layers of a star while the inner core remains intact (partial TDEs). This is the case when the relative impact parameter $\beta=r_{\rm t}/r_{\rm p}<\beta_{\rm crit}$, where $r_{\rm t}$ is the tidal radius and $r_{\rm p}$ is the pericenter distance of the star. For main-sequence stars, $\beta_{\rm crit}$ was constrained to be $\sim 2-3$[140] and for these values, the star likely gets fully disrupted. Partial TDEs exhibit variability in both UV/optical and X-ray wavebands, often with a time delay due to different processes (e.g. AT2018fyk[141] or ASASSN0-14ko[142]).
Currently the shortest repeating nuclear transients associated with galactic nuclei are Quasiperiodic Erupters (QPEs [143, 144]), which exhibit large-amplitude soft X-ray flares on the timescale of hours to days; however, they do not exhibit significant variability in the optical/UV domains; see, however, the analysis of the “Ansky” source that shows signs of UV variability[23]. Recently, a class of repeating soft X-ray nuclear transients on the intermediate timescales of tens of days has been revealed (represented by Swift J0230-28[20, 21]), which shares most of the spectral properties with QPEs, and they likely correspond to partial TDEs of non-main-sequence stars, such as red giants or even gas-giant planets[22]. In the case of Swift J0230-28, no significant variability in the optical/UV bands was found, though a radio flare was associated with the start of one of the X-ray flares[21].
@@
The peak soft X-ray luminosity of QPE outbursts reaches $\sim 10^{42}$ erg s-1. The exact mechanism for their production is uncertain – the proposed models span from accretion-disk intrinsic instabilities[138, 145] to bodies periodically crossing an accretion disk[146, 139]. A favored class of models involve extreme mass-ratio inspirals (EMRIs)[147], i.e. stars or compact remnants which collide with the accretion disks of supermassive black holes (SMBHs), producing flares through disk–star interactions involving shocks within the ejected material [148]. This class of models can also address the variability in the flare recurrence timescale, in particular the long-short recurrence timescale pattern, which can be attributed to the EMRI prograde orbital (Schwarzschild) precession in combination with the accretion-disk precession [148, 149].
@@
Since the EMRI distance can be expressed in gravitational radii as $r_{\rm EMRI}/r_{\rm g}\propto M_{\bullet}^{-2/3}P_{\rm QPE}^{2/3}$, longer-period QPEs are expected to result in softer flares peaking in the UV domain for the fixed SMBH mass[150]. Similarly, for the QPE periodicity comparable to the X-ray QPEs, the flares are expected to be softer and potentially detectable in the UV domain in lower-mass SMBH systems. Also, since some X-ray QPEs seem to be associated with post-TDE nuclei with a decreasing relative accretion rate, the flare spectra are expected to get softer with the drop in the accretion rate, potentially ceasing to be detected in the X-ray domain and appearing in the UV bands [150] when the relative accretion rate reaches $\dot{m}\sim\mathcal{O}(10^{-2})$. The detection of such events in the UV domain is feasible with future missions, such as ULTRASAT and QUVIK, which have sensitivities down to $\sim 22$ AB magnitude. The QUVIK mission has the projected sensitivity limit of $\sim 22$ AB mag in the NUV band (325 nm), which implies the monochromatic flux density limit of $F_{\nu}\sim 5.75\times 10^{-29}\,{\rm erg\,s^{-1}\,cm^{-2}\,Hz^{-1}}$ or $\nu F_{\nu}\sim 5.31\times 10^{-14}\,{\rm erg\,s^{-1}\,cm^{-2}}$ in the NUV band. UV QPEs in galaxies could thus be detectable within a distance of up to $d_{\rm L}=235(L_{\rm QPE}/3.5\times 10^{41}\,{\rm erg\,s^{-1}})^{1/2}(\nu F_{\nu}/5.31\times 10^{-14}\,{\rm erg\,s^{-1}\,cm^{-2}})^{-1/2}\,{\rm Mpc}$ ($z=0.053$), assuming the typical UV QPE luminosity of $3.5\times 10^{41}\,{\rm erg\,s^{-1}}$[150]. The detection and monitoring of UV QPEs would significantly advance our understanding of repeating nuclear transients, their potential association with EMRIs [147, 139], and the evolution of SMBH accretion disks plausibly following a TDE[151].
QUVIK observational strategy: With the high-cadence monitoring of selected candidates, QUVIK can detect quasiperiodicity at the level of $\sim 0.2$ days or about 5 hours. This way, the repeating UV nuclear transients can be extended towards significantly shorter periods. It will be necessary to conduct longer monitoring for a few months so that the periodic signal can effectively be disentangled from stochastic variability. The nature of the suggested UV variability of QPE sources[23] can thus be better constrained. In combination with the X-ray data, it will be possible to test the origin of UV flares, specifically if they originate due to the reprocessing of harder radiation or if they are produced in shocks due to star-disk interactions or alternatively due to accretion-disk instabilities. The sample of detectable repeating nuclear UV transients is located towards lower redshifts. When we adopt the redshift distribution of detected 13 X-ray QPEs [139], we obtain the median redshift of $z=0.024$ with the 16% and the 84% percentiles of $0.018$ and $0.091$, respectively (the mean redshift is $0.047\pm 0.036$, while the minimum and the maximum redshifts are 0.0151 and 0.1177, respectively). For the peak UV luminosity of $L_{\rm 300}\sim 10^{42}\,{\rm erg\,s^{-1}}$ at 300 nm (NUV band)[150], we obtain the AB magnitude $m_{\rm AB}=-2.5\log{F_{\nu}}-48.60\sim 22.4$ for $z=0.1$. Taking the conservative sensitivity limit of $m_{\rm AB}\lesssim 21.5$, we infer the redshift limit of $z\lesssim 0.065$ for the detection of QPE-like UV transients, which is consistent with the estimate above considering the uncertainties of peak UV luminosities and of the sensitivity limit. A new class of (repeating) UV transients can also be detected in connection with blazar jets and related $\gamma$-ray flares[152]. Interactions of stars orbiting the SMBH with the nuclear outflow and the jet could result in recurring shocks[153, 154, 155, 156] with the emission spectrum peaking at progressively lower energies as the shocked material cools downstream. The UV counterparts of the blazar sources could be searched for within the dedicated observational program aimed at radio-loud AGN sources.
5 Synergy of QUVIK observations with other missions
@@
    the telescope will perform dedicated monitoring of the variability of selected bright AGN ($m_{\rm AB}\lesssim 18$ mag) located at low to intermediate redshifts ($z\lesssim 1$). These AGN programs and campaigns will be selected based on a competitive peer-review process following annual proposal calls available to the international community. After the established proprietary period, the data will be freely available via the archive portal.
@@ 6 Summary
QUVIK is an approved versatile two-band UV photometry space mission currently under development. It is designed to perform observations of the early UV emission of kilonovae, which is its primary scientific objective. In addition, it will conduct monitoring and follow-up observations of various sources, including stellar transients, hot stars, binary stars, stellar associations and clusters. Building on the legacy of previous successful UV space telescopes, QUVIK will also bring valuable results concerning active galactic nuclei (AGN) and nuclear transients. The main areas of AGN research include:
  • reverberation mapping of accretion disks. QUVIK can conduct high-cadence observations of AGN variability simultaneously in both FUV and NUV bands. For the sources with the SMBH mass of $10^{9}\,M_{\odot}$, the required cadence is $\lesssim 1$ day, while for $10^{8}\,M_{\odot}$, the cadence needs to increase to $\sim 0.1$ days, while the campaign duration should last for about half a year,
  • the monitoring duration will need to be adjusted taking into account the redshift of the nuclear source; however, because of the sensitivity limits in the NUV and the FUV bands, the focus will be on relatively low-redshift sources with $z\lesssim 0.5$, for which the time dilation factor is $(1+z)\lesssim 1.5$. The cadence will remain to be $\lesssim 1$ day since for higher-redshift sources the UV telescope probes progressively shorter-wavelength spectral regions that vary on shorter timescales in the rest frame,
@@
QUVIK will operate in synergy with wide field-of-view observing facilities, including ULTRASAT, Vera C. Rubin Legacy Survey of Space and Time as well as with gravitational-wave observatories (LIGO-Virgo-KAGRA). QUVIK, with its agile mode of operation and fast response, will also provide important synergies for the larger future UVEX mission equipped with a 2-degree-long, multi-width UV spectrometer across a wide spectral band (115-265 nm, $R>1000$) in addition to a two-band FUV-NUV imager [58]. Importantly, its potential detections of repeating nuclear transients, some of which may be associated with extreme- and intermediate-mass ratio inspirals, will be beneficial for the upcoming space-borne low-frequency gravitational-wave observatories (LISA) in terms of constraining their cosmological volume densities and looking for potential electromagnetic counterparts of gravitational-wave sources [147].
Data availability statement
All the data used and obtained in this manuscript will be made available upon request.
Disclosure statement
The authors declare there are no financial interests, commercial affiliations, or other potential conflicts of interest that have influenced the objectivity of this research or the writing of this paper.
@@ Acknowledgments
MZ, HB, MP, VK, MM, and PS acknowledge the support from the Czech Science Foundation (GACR) Junior Star grant no. GM24-10599M. NW, LŠ, and ILG were supported by the Czech Science Foundation (GACR) grant GX21-13491X. PK and ML acknowledge the continuing support of the OPUS-LAP/GAČR-LA bilateral research grant (2021/43/I/ST9/01352/OPUS22 and GF23-04053L). JL acknowledges the support from the J.W. Fulbright Commission in the Czech Republic. JŘ acknowledges the support from the Czech Science Foundation (GAČR) grant no. GC24-11487J.
Biographies and photographs of the other authors are not available.